L=(3x+20) m=(2X+40)After the problem is solved, they both have to be equal. Can you help me with this?
Answer by checkley77(12569)

L=(3x+20) m=(2X+40)
If you mean L=m then:
3x+20=2x+40
3x-2x=40-20
x=20 Answer.
Proof:
3*20+20=2*20+40
60+20=40+40
80=80